Determine the boiling point of a solution of 10 grams of a substance of molar mass of 94g/mol in 500 grams of water


1
Expert's answer
2021-06-01T02:52:20-0400

Mass of substance"(w)=10g"

Molecular weight of substance"(Gmw)=95g\/mol"

Volume of "H_2O(v)=500gms=500ml"

Molarity of solution"(M)=" "W\over Gmw" "\times" "1000\over v(ml)"

"M=" "10\over 94" "\times" "1000\over 500" "=0.2127M"

Change in boiling point ΔTb​=Kb​×M

"\Delta T_b=0.512°c/_m\times 0.2127M"

"\Delta T_b=0.1089°C"

Boiling point"=100\u00b0C+0.1089\u00b0C=100.1089\u00b0C"